# Towards a universal QAOA protocol: Evidence of quantum advantage in solving combinatorial optimization problems

### v1

@@ III.3 Experiments
Figure 9: (a) 109-qubit W-MaxCut experiment on IBM Eagle devices using LR-QAOA. The upper-left plot represents the best approximation ratio observed from the 10000 samples at each $p$. The bottom-left plot shows the average approximation ratio where the dotted line represents the raw result from ibm_brisbane and the solid lines represent the mitigated results over the three different IBM devices. The black dashed line is the limit where the system reaches the maximally mixed state. The right plot shows the IBM Eagle layout with the 1D random W-MaxCut on it. The colors represent the random weights chosen from 4 possible values 0.1, 0.2, 0.5, and 2.0 for each edge in the graph. The inset shows how the 2-qubit gates are implemented using only a depth of 2 for each LR-QAOA layer. (b) 25-qubit W-MaxCut experiment comparative analysis for 3 different vendors. The bottom-left plot presents the graph of the W-MaxCut selected with the corresponding weight values. In the upper plot, the approximation ratio of the devices with the triangles, circles, diamonds, and pentagons correspond to the mitigated results from a noiseless simulation, ibm_brisbane, ionq_aria, and quantinuum_H2, respectively. Because of a limitation in the maximum number of single-qubit and two-qubit gates, there are no results above $p=15$ using ionq_aria. In the bottom left plot the probability of success for the same experiment.
@@ IV Conclusions
In this work, we have presented analytical, numerical, and experimental evidence that LR-QAOA constitutes a universal QAOA protocol. This means that this protocol works efficiently independently of the problem tested, increasing the probability of success as the number of layers increases. We simulate MIS, BPP, TSP, MaxCut, W-MaxCut, 3-MaxCut, KP, PO, Max-2-SAT, and Max-3-SAT problems with up to 42 qubits and 400 layers on the modular supercomputer JUWELS [20]. Additionally, we test LR-QAOA using W-MaxCut problems from 5 to 109-qubit cases and $p$ from 3 to 100 on real quantum hardware using ibm_brisbane, ibm_osaka, ibm_kyoto, quantinuum_H2, and ionq_aria finding that LR-QAOA is resilient to noise. We show that this characteristic comes from the algorithm’s ability to overcome noise. While noise pushes the system through a maximal mixed state, LR-QAOA pushes to a minimum energy correcting errors on the way.
@@
We observe that the success probability of the optimal solutions using LR-QAOA for the different COPs seems to scale as $probability(x^{*})\approx 1/2^{(\eta N_{q}/p)}$ for a constant $\eta$. We add further evidence that this is the case using W-MaxCut, MaxCut, Max-2-SAT, and Max-3-SAT with up to 42 qubits. Using $p=N_{q}$, we find that the $probability(x^{*})=\mathrm{cte}$ remains nearly constant for W-MaxCut, MaxCut, and Max-3-SAT. The Max-2-SAT case is an exception, using $p=N_{q}$, it already shows an exponential decay in the probability of success. We think this is a consequence of problems with a high concentration of solutions close to the optimal solution. With all of this information, we conjecture that $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for general problems.
@@
We compare SA, LR-QAOA, and CPL EX in terms of TTS for fully connected W-MaxCut problems. In the largest case, $N_{q}=42$, LR-QAOA has an improvement of 2 orders of magnitude compared to the other two solvers. We think there is a fundamental improvement in LR-QAOA. While any other classical solver needs time to learn the energy structure of the problem, LR-QAOA exploits the $U_{C}$ phenomenon, Eq. 7, having the full energy information stored in the rotation of the amplitudes. This information is then efficiently used by the $U_{B}$ gate to amplify the probability of obtaining low-energy solutions.
Moreover, we find that LR-QAOA is resilient to noise. This is important as we are at a stage where quantum computers have moderate noise. We simulate a MIS using LR-QAOA with $p=80$ and at $p=15$ we add a layer of X gates to see how the algorithm evolves under this noise. We find that there is an interplay between noise and the algorithm trying to get to the ground state. If the noise is low enough, the own algorithm corrects it to some extent. We use depolarizing noise for the simulation of a 10-qubit W-MaxCut problem using the LR-QAOA protocol and find that it fits well with the experimental results on ionq_aria.
@@
The 109-qubit experiment already in the regime of quantum supremacy constitutes indirect proof that the LR-QAOA protocol works for larger cases. These cases constitute a set of experiments that cannot be simulated exactly on classical computers, and even approximations will result hard as $p$ grows. We find that some information on the LR-QAOA algorithm remains for the largest case, $N_{q}=109$ and $p=100$.

### v2

@@ III.3 Experiments
Figure 9: (a) 109-qubit W-MaxCut experiment on IBM Eagle devices using LR-QAOA. The upper-left plot represents the best approximation ratio observed from the 10000 samples at each $p$. The bottom-left plot shows the average approximation ratio where the dotted line represents the raw result from ibm_brisbane and the solid lines represent the mitigated results over the three different IBM devices. The black dashed line is the limit where the system reaches the maximally mixed state. The right plot shows the IBM Eagle layout with the 1D random W-MaxCut on it. The colors represent the random weights chosen from 4 possible values 0.1, 0.2, 0.5, and 2.0 for each edge in the graph. The inset shows how the 2-qubit gates are implemented using only a depth of 2 for each LR-QAOA layer. (b) 25-qubit W-MaxCut experiment comparative analysis for 3 different vendors. The bottom-left plot presents the graph of the W-MaxCut selected with the corresponding weight values. In the upper plot, the approximation ratio of the devices with the triangles, circles, diamonds, and pentagons correspond to the mitigated results from a noiseless simulation, ibm_brisbane, ionq_aria, and quantinuum_H2, respectively. Because of a limitation in the maximum number of single-qubit and two-qubit gates, there are no results above $p=15$ using ionq_aria. In the bottom right plot the probability of success for the same experiment.
@@ IV Conclusions
In this work, we have presented numerical and experimental evidence that LR-QAOA constitutes a universal QAOA protocol. This means that this protocol works efficiently independently of the problem tested, increasing the probability of success as the number of layers increases. We simulate MIS, BPP, TSP, MaxCut, W-MaxCut, 3-MaxCut, KP, PO, Max-2-SAT, and Max-3-SAT problems with up to 42 qubits and 400 layers on the modular supercomputer JUWELS. Additionally, we test LR-QAOA using W-MaxCut problems from 5 to 109-qubit cases and $p$ from 3 to 100 on real quantum hardware using ibm_brisbane, ibm_osaka, ibm_kyoto, quantinuum_H2, and ionq_aria finding that LR-QAOA is resilient to noise. We show that this characteristic comes from the algorithm’s ability to overcome noise. While noise pushes the system through a maximal mixed state, LR-QAOA pushes to a minimum energy correcting errors on the way.
@@
We observe that the success probability of the optimal solutions using LR-QAOA for the different COPs seems to scale as $probability(x^{*})\approx 1/2^{(\eta N_{q}/p)}$ for a constant $\eta$. We add further evidence that this is the case using W-MaxCut, MaxCut, Max-2-SAT, and Max-3-SAT with up to 42 qubits. Using $p=N_{q}$, we find that on average, the $probability(x^{*})$ remains nearly constant for W-MaxCut, MaxCut, and Max-3-SAT. The Max-2-SAT case is an exception, using $p=N_{q}$, it already shows an exponential decay in the probability of success. We think this is a consequence of problems with a high concentration of solutions close to the optimal solution. With all of this information, we conjecture that $probability(x^{*})=1/2^{(\eta N_{q}/p)}$ for general problems.
@@
We compare SA, LR-QAOA, and CPLEX in terms of TTS for fully connected W-MaxCut problems, observing a scaling improvement in LR-QAOA. We attribute this advantage to the unique characteristics of LR-QAOA. While classical solvers require time to learn the energy structure of the problem, LR-QAOA leverages the $U_{C}$ phenomenon (Eq. 7), which inherently contains the full energy information in the rotation of the amplitudes. This information is then efficiently utilized by the $U_{B}$ gate to amplify the probability of obtaining low-energy solutions.
Moreover, we find that LR-QAOA is resilient to noise. This is important as we are at a stage where quantum computers have moderate noise. We simulate a MIS using LR-QAOA with $p=80$ and at $p=15$ we add a layer of X gates to see how the algorithm evolves under this noise. We find that there is an interplay between noise and the algorithm trying to get to the ground state. If the noise is low enough, the algorithm corrects it to some extent. We use depolarizing noise for the simulation of a 10-qubit W-MaxCut problem using the LR-QAOA protocol and find that it fits well with the experimental results on ionq_aria.
@@
The 109-qubit experiment already in the regime of quantum supremacy constitutes indirect proof that the LR-QAOA protocol works for larger cases. These cases constitute a set of experiments that cannot be simulated exactly on classical computers, and even approximations will become hard as $p$ grows. We find that some information on the LR-QAOA algorithm remains for the largest case, $N_{q}=109$ and $p=100$.
@@ A.1 Max-3-SAT problem
A.2 2D Performance Diagram
Figure 10 shows the success probability landscape of LR-QAOA for values of $\Delta_{\beta}$ and $\Delta_{\gamma}$ ranging from 0 to $43\pi/$ 4 at $p=50$ across different COPs. Surprisingly, consistent behavior is observed across the different plots, although the structure of the Hamiltonian varies considerably between them. As discussed in Sec. II.6, such consistency is only observed when the Hamiltonian is normalized. The red region in these plots resembles the ridge region of the performance diagram, indicating that excessively large values of $\Delta_{\beta}$ or $\Delta_{\gamma}$ do not lead to improvement. Moreover, the best performance does not necessarily occur when $\Delta_{\beta}=\Delta_{\gamma}$, suggesting potential performance improvement through parameter tuning. In our case, a $\Delta_{\beta}\leq 0.6$ and a $\Delta_{\gamma}\leq 0.6$ give favorable results.
Figure 10: 2D Performance diagram of $\Delta_{\beta}$ vs. $\Delta_{\gamma}$ for the LR-QAOA protocol with p=50 for a random instance of (a) 12-qubit MIS, (b) 12-qubit KP, (c) 9-qubit TSP, (d) 12-qubit PO, (e) 12-qubit BPP, and (f) a 16-qubit W-MaxCut.